224 man-hours are required to complete a project. 4 men are employed at $7.50 per hour for the job. The normal working hours for the men are from 0900h to 1800h with a one-hour lunch break in between. As hourly-rated workers, they are not paid during the lunch hour.
(i) How much would each man be paid for completing the project?
(ii) Under the Labour Law, the workers are entitled to overtime pay of 1.5 times their hourly rate when they work beyond the normal working hours. The project became urgent and needs to be completed in 4 days. How much would the total labour cost be now?
(iii) What is the percentage increase in labour cost due to the early completion of the project?

Answer by robertb(5830)

(i) Each man would have to work 224/4 = 56 hours for the entire project. This means he would be paid (8 hours/day)*($7.50/hour)*(7 days) = $420. (Project is completed in 7 days.)
(ii) Any hours in excess of 8 hours is paid at a rate of 1.5*$7.50 = $11.25. Since the project has to be finished in 4 days, he has to work 56/4 = 14 hours per day, the first 8 of which is paid at $7.50 per hour, and the last 6 of which is paid at $11.25 per hour.
Hence, the total pay for each man is
(8 hours/day)*($7.50/hour)*(4 days) + (6 hours/day)*($11.25/hour)*(4 days) = $510.
(iii)The percentage increase in labor cost due to early completion would be ($510 - $420)/$420 *100% = 21.43%.